### v1

# Software paper for submission to the Journal of Open Research Software

Abstract
PDSim is an R package that enables users to simulate commodity futures prices using the polynomial diffusion model introduced in Filipović and Larsson [7] through both a Shiny web application and R scripts. It also provides state variables and contract estimations via the Extended Kalman Filter (EKF) or Unscented Kalman Filter (UKF). With its user-friendly interface, PDSim makes the features of simulations and estimations accessible to all users. To date, it is the only package specifically designed for the simulation and estimation of the polynomial diffusion model. Additionally, the package integrates the Schwartz and Smith two-factor model [17] as an alternative approach. PDSim offers versatile deployment options, including running locally, via the Shiny server, or through Docker.
Keywords
Polynomial diffusion model; Schwartz-Smith model; commodity futures; two-factor modelling
Introduction
Stochastic models are crucial in the analysis of commodity futures, serving important roles in various financial areas such as price forecasting, risk management, and portfolio optimisation. Typically, the underlying spot price, denoted as $S_{t}$, is modelled as a function of certain factors. In 1990, the Ornstein-Uhlenbeck (OU) process was introduced to model oil futures in a two-factor setup, representing spot price and convenience yield [8]. Building on this foundation, Schwartz and Smith [17] modelled the logarithm of the underlying spot price of crude oil futures as the sum of two hidden factors. These factors, assumed to follow the OU process, capture short-term fluctuations and the long-term equilibrium price level, respectively. This latent factor model and its extensions have been widely used in stochastic modelling in various fields, including biology [6], electricity forwards [13], agricultural commodity futures [18], and crude oil futures [2, 5, 16].
Despite its widespread use, this framework has certain limitations. First, a closed-form expression is required for futures prices. To achieve this, it is common to assume that all involved factors follow a Gaussian distribution and that the logarithm of the spot price is a linear function of these factors. This assumption ensures a log-normal distribution of the spot price. Second, within this framework, the spot price is typically modelled in logarithmic form, which generally works well. However, an unprecedented event on April 20, 2020, challenged this approach: the front-month May 2020 WTI crude oil futures settled at an extraordinary value of -\$37.63 per barrel on the New York Mercantile Exchange. This rare phenomenon significantly undermines the validity of the entire framework.
The polynomial diffusion model was introduced to overcome these limitations. In this framework, the spot price is represented as a polynomial of any order in terms of the factors. In particular, under certain conditions, it can be proven that the conditional expectation of the spot price, which is equivalent to the futures price under the assumption of an arbitrage-free market and non-stochastic interest rate, is also a polynomial in terms of factors. The mathematical foundations of the polynomial diffusion model were introduced in [7], with applications of this framework seen in the modelling of electricity forwards, where the spot price is represented by a quadratic form of two factors [14].
I. Schwartz and Smith two-factor model
Under the Schwartz-Smith framework, the logarithm of spot price $S_{t}$ is modelled as the sum of two factors $\chi_{t}$ and $\xi_{t}$,
$\log{(S_{t})}=\chi_{t}+\xi_{t},$ (1)
where $\chi_{t}$ represents the short-term fluctuation and $\xi_{t}$ is the long-term equilibrium price level. Additionally, we assume both $\chi_{t}$ and $\xi_{t}$ follow a risk-neutral Ornstein-Uhlenbeck process,
$d\chi_{t}=(-\kappa\chi_{t}-\lambda_{\chi})dt+\sigma_{\chi}dW_{t}^{\chi*},$ (2)
and
$d\xi_{t}=(\mu_{\xi}-\gamma\xi_{t}-\lambda_{\xi})dt+\sigma_{\xi}dW_{t}^{\xi*},$ (3)
where $\kappa,\gamma\in\mathbb{R}^{+}$ are the speed of mean-reversion parameters, $\mu_{\xi}\in\mathbb{R}$ is the mean level of the long-term factor, $\sigma_{\chi},\sigma_{\xi}\in\mathbb{R}^{+}$ are the volatility parameters, and $\lambda_{\chi},\lambda_{\xi}\in\mathbb{R}$ are risk premiums. The processes $(W_{t}^{\chi*})_{t\geq 0}$ and $(W_{t}^{\xi*})_{t\geq 0}$ are correlated standard Brownian Motions with
$\mathbb{E}\left(dW_{t}^{\chi*}dW_{t}^{\xi*}\right)=\rho dt.$
In the original Schwartz-Smith model [17], the long-term factor $\xi_{t}$ is geometric Brownian motion (gBm), corresponding to setting $\gamma=0$. Since 2000, many studies such as [2, 4, 15, 16] and others considered the extended version of the Schwartz-Smith model by introducing $\gamma\geq 0$. For simplicity, further, we continue to refer to this model with $\gamma\geq 0$ as the Schwartz-Smith model. Setting $\lambda_{\chi}=\lambda_{\xi}=0$ in Equation 2 and Equation 3 gives the real processes of the factors. The risk-neutral processes are used to calculate futures prices, and the real processes are for modelling state variables in real time.
In discrete time, $\chi_{t}$ and $\xi_{t}$ are jointly normally distributed. Therefore, the spot price is log-normally distributed. Moreover, under the arbitrage-free assumption and non-stochastic interest rate, the futures price $F_{t,T}$ at current time $t$ with maturity time $T$ must be equal to the expected value of spot price at maturity time $T$,
$F_{t,T}=\mathbb{E}^{*}(S_{T}|\mathcal{F}_{t}),$
where $\mathcal{F}_{t}$ is the information known at time $t$ and $\mathbb{E}^{*}(\cdot)$ is the expectation under the risk-neutral processes from Equation 2 and Equation 3. Then we can get the linear Gaussian state space model:
$x_{t}=c+Ex_{t-1}+w_{t},$ (4)
$y_{t}=d_{t}+F_{t}x_{t}+v_{t},$ (5)
where $x_{t}=\left[\begin{matrix}\chi_{t}\\ \xi_{t}\end{matrix}\right],c=\left[\begin{matrix}0\\ \frac{\mu_{\xi}}{\gamma}\left(1-e^{-\gamma\Delta t}\right)\end{matrix}\right],E=\left[\begin{matrix}e^{-\kappa\Delta t}&0\\ 0&e^{-\gamma\Delta t}\end{matrix}\right],y_{t}=\left(\log{(F_{t,T_{1}})},\dots,\log{(F_{t,T_{m}})}\right)^{\top}$, $d_{t}=\left(A(T_{1}-t),\dots,A(T_{m}-t)\right)^{\top},F_{t}=\left[\begin{matrix}e^{-\kappa(T_{1}-t)},\dots,e^{-\kappa(T_{m}-t)}\\ e^{-\gamma(T_{1}-t)},\dots,e^{-\gamma(T_{m}-t)}\end{matrix}\right]^{\top}$ and $m$ is the number of futures contracts. The function $A(\cdot)$ is given by
$A(\tau)=$ $-\frac{\lambda_{\chi}}{\kappa}(1-e^{-\kappa\tau})+\frac{\mu_{\xi}-\lambda_{\xi}}{\gamma}(1-e^{-\gamma\tau})$
$+\frac{1}{2}\left(\frac{1-e^{-2\kappa\tau}}{2\kappa}\sigma_{\chi}^{2}+\frac{1-e^{-2\gamma\tau}}{2\gamma}\sigma_{\xi}^{2}+2\frac{1-e^{-(\kappa+\gamma)\tau}}{\kappa+\gamma}\sigma_{\chi}\sigma_{\xi}\rho\right),$
where $\tau$ is the time to maturity. $w_{t}$ and $v_{t}$ are multivariate Gaussian noises with mean 0 and covariance matrix $\Sigma_{w}$ and $\Sigma_{v}$ respectively, where
$\Sigma_{w}=\left[\begin{matrix}\frac{1-e^{-2\kappa\Delta t}}{2\kappa}\sigma_{\chi}^{2}&\frac{1-e^{-(\kappa+\gamma)\Delta t}}{\kappa+\gamma}\sigma_{\chi}\sigma_{\xi}\rho\\ \frac{1-e^{-(\kappa+\gamma)\Delta t}}{\kappa+\gamma}\sigma_{\chi}\sigma_{\xi}\rho&\frac{1-e^{-2\gamma\Delta t}}{2\gamma}\sigma_{\xi}^{2}\end{matrix}\right],$
and we assume $\Sigma_{v}$ is diagonal, i.e.,
$\Sigma_{v}=\left[\begin{matrix}\sigma_{1}^{2}&0&\dots&0\\ 0&\sigma_{2}^{2}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\sigma_{m}^{2}\end{matrix}\right].$
$\Delta t$ is the time step. Under this framework, $c,E,\Sigma_{w}$ and $\Sigma_{v}$ are deterministic but $d_{t}$ and $F_{t}$ are time-variant.
II. Polynomial diffusion model
In this section, we present a general framework of the polynomial diffusion model. A simulation study is given in [10].
Under the polynomial diffusion framework, the spot price $S_{t}$ is expressed as a polynomial function of the hidden state vector $x_{t}$ (with components $\chi_{t}$ and $\xi_{t}$):
$S_{t}=p_{n}(x_{t})=\alpha_{1}+\alpha_{2}\chi_{t}+\alpha_{3}\xi_{t}+\alpha_{4}\chi_{t}^{2}+\alpha_{5}\chi_{t}\xi_{t}+\alpha_{6}\xi_{t}^{2}.$
In this context, we assume that the polynomial $p_{n}(x_{t})$ has a degree of 2 (or 1 if $\alpha_{4}=\alpha_{5}=\alpha_{6}=0$). However, it is worth noting that all the theorems presented here are applicable even for polynomials of higher degrees. Additionally, similar to the Schwartz and Smith model, we assume that $\chi_{t}$ and $\xi_{t}$ follow an Ornstein–Uhlenbeck process
$d\chi_{t}=-\kappa\chi_{t}dt+\sigma_{\chi}dW_{t}^{\chi}$
$d\xi_{t}=(\mu_{\xi}-\gamma\xi_{t})dt+\sigma_{\xi}dW_{t}^{\xi}$
for real processes and
$d\chi_{t}=(-\kappa\chi_{t}-\lambda_{\chi})dt+\sigma_{\chi}dW_{t}^{\chi*}$
$d\xi_{t}=(\mu_{\xi}-\gamma\xi_{t}-\lambda_{\xi})dt+\sigma_{\xi}dW_{t}^{\xi*}$
for risk-neutral processes.
Now, consider any processes that follow the stochastic differential equation
$dX_{t}=b(X_{t})dt+\sigma(X_{t})dW_{t},$
where $W_{t}$ is a $d$-dimensional standard Brownian Motion and map $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ is continuous. Define $a:=\sigma\sigma^{\top}$. For maps $a:\mathbb{R}^{d}\to\mathbb{S}^{d}$ and $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$, suppose we have $a_{ij}\in Pol_{2}$ and $b_{i}\in Pol_{1}$. $\mathbb{S}^{d}$ is the set of all real symmetric $d\times d$ matrices and $Pol_{n}$ is the set of all polynomials of degree at most $n$. Then the solution of the SDE is a polynomial diffusion. Moreover, we define the generator $\mathcal{G}$ associated to the polynomial diffusion $X_{t}$ as
$\mathcal{G}f(x)=\frac{1}{2}Tr\left(a(x)\nabla^{2}f(x)\right)+b(x)^{\top}\nabla f(x)$
for $x\in\mathbb{R}^{d}$ and any $C^{2}$ function $f$. Let $N$ be the dimension of $Pol_{n}$, and $H:\mathbb{R}^{d}\to\mathbb{R}^{N}$ be a function whose components form a basis of $Pol_{n}$. Then for any $p\in Pol_{n}$, there exists a unique vector $\vec{p}\in\mathbb{R}^{N}$ such that
$p(x)=H(x)^{\top}\vec{p}$
and $\vec{p}$ is the coordinate representation of $p(x)$. Moreover, there exists a unique matrix representation $G\in\mathbb{R}^{N\times N}$ of the generator $\mathcal{G}$, such that $G\vec{p}$ is the coordinate vector of $\mathcal{G}p$. So we have
$\mathcal{G}p(x)=H(x)^{\top}G\vec{p}.$
Theorem 1: Let $p(x)\in Pol_{n}$ be a polynomial with coordinate representation $\vec{p}\in\mathbb{R}^{N}$, $G\in\mathbb{R}^{N\times N}$ be a matrix representation of generator $\mathcal{G}$, and $X_{t}\in\mathbb{R}^{d}$ satisfies the SDE. Then for $0\leq t\leq T$, we have
$\mathbb{E}\left[p(X_{T})|\mathcal{F}_{t}\right]=H(X_{t})^{\top}e^{(T-t)G}\vec{p},$
where $\mathcal{F}_{t}$ represents all information available until time $t$.
Obviously, the hidden state vector $x_{t}$ satisfies the SDE with
$b(x_{t})=\left[\begin{matrix}-\kappa\chi_{t}-\lambda_{\chi}\\ \mu_{\xi}-\gamma\xi_{t}-\lambda_{\xi}\end{matrix}\right],\sigma(x_{t})=\left[\begin{matrix}\sigma_{\chi}&0\\ 0&\sigma_{\xi}\end{matrix}\right],a(x_{t})=\sigma(x_{t})\sigma(x_{t})^{\top}=\left[\begin{matrix}\sigma_{\chi}^{2}&0\\ 0&\sigma_{\xi}^{2}\end{matrix}\right].$
The basis
$H(x_{t})=(1,\chi_{t},\xi_{t},\chi_{t}^{2},\chi_{t}\xi_{t},\xi_{t}^{2})^{\top}$
has a dimension $N=6$. The coordinate representation is
$\vec{p}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{6})^{\top}.$
By applying $\mathcal{G}$ to each element of $H(x_{t})$, we get the matrix representation
$G=\left[\begin{matrix}0&-\lambda_{\chi}&\mu_{\xi}-\lambda_{\xi}&\sigma_{\chi}^{2}&0&\sigma_{\xi}^{2}\\ 0&-\kappa&0&-2\lambda_{\chi}&\mu_{\xi}-\lambda_{\xi}&0\\ 0&0&-\gamma&0&-\lambda_{\chi}&2\mu_{\xi}-2\lambda_{\xi}\\ 0&0&0&-2\kappa&0&0\\ 0&0&0&0&-\kappa-\gamma&0\\ 0&0&0&0&0&-2\gamma\end{matrix}\right].$
Then, by Theorem 1, the futures price $F_{t,T}$ is given by
$F_{t,T}=\mathbb{E}^{*}(S_{T}|\mathcal{F}_{t})=H(x_{t})^{\top}e^{(T-t)G}\vec{p}.$
Therefore, we have the non-linear state-space model
$x_{t}=c+Ex_{t-1}+w_{t},$
$y_{t}=H(x_{t})^{\top}e^{(T-t)G}\vec{p}+v_{t}.$
1 III. Filtering methods
Figure 1: Flowchat of EKF.
Figure 2: Flowchat of UKF.
The Kalman Filter (KF) [9] is a commonly used filtering method in estimating hidden state variables. However, KF can only deal with the linear Gaussian state model. To capture the non-linear dynamics in the PD model, we use Extended Kalman Filter (EKF) [11] and Unscented Kalman Filter (UKF) [12, 19]. Suppose we have a non-linear state-space model
$x_{t}=f(x_{t-1})+w_{t},w_{t}\sim N(\textbf{0},\Sigma_{w}),$
$y_{t}=h(x_{t})+v_{t},v_{t}\sim N(\textbf{0},\Sigma_{v}).$
The EKF linearises the state and measurement equations through the first-order Taylor series. To run KF, we replace $J_{f}$ and $J_{h}$ with $E$ and $F_{t}$ respectively, where $J_{f}$ and $J_{h}$ are the Jacobians of $f(\cdot)$ and $h(\cdot)$. In contrast, the UKF uses a set of carefully chosen points, called sigma points, to represent the true distributions of state variables. Then, these sigma points are propagated through the state equation. The flowcharts of EKF and UKF are given in Figure 1 and Figure 2, where some notations are defined as follow:
$a_{t|t-1}$ $:=\mathbb{E}(x_{t}|\mathcal{F}_{t-1}),$ $P_{t|t-1}$ $:=Cov(x_{t}|\mathcal{F}_{t-1}),$
$a_{t}$ $:=\mathbb{E}(x_{t}|\mathcal{F}_{t}),$ $P_{t}$ $:=Cov(x_{t}|\mathcal{F}_{t}).$
In this application, we use KF for the Schwartz-Smith model, and EKF/UKF for the polynomial diffusion model.
IV. Comparison with existing libraries
The R package “NFCP” [3] was developed for multi-factor pricing of commodity futures, which is a generalisation of the Schwartz-Smith model. However, this package doesn’t accommodate the polynomial diffusion model. There are no R packages available for PD models currently.
There are many packages in R for Kalman Filter, for example, “dse”, “FKF”, “sspir”, “dlm”, “KFAS”: “dse” can only take time-invariant state and measurement transition matrices; “FKF” emphasizes computation speed but cannot run smoother; “sspir”, “dlm” and “KFAS” have no deterministic inputs in state and measurement equations. For the non-linear state-space model, the functions “ukf” and “ekf” in package “bssm” run the EKF and UKF respectively. However, this package was designed for Bayesian inference where a prior distribution of unknown parameters is required. To achieve the best collaboration of filters and models, we developed functions of KF, EKF and UKF within this code.
Implementation and architecture
The graphical user interface (GUI) provides an easy way for anyone to use the PDSim package, even without programming knowledge. Simply enter the necessary parameters, and PDSim will simulate the data and generate well-designed interactive visualizations. Currently, PDSim supports data simulation from two models: the Schwartz and Smith two-factor model [17] and the polynomial diffusion model [7]. In this section, we will guide you through using the GUI to simulate data.
I. Implementation of the Schwartz-Smith model
Figure 3: Establish global configurations.
First, we establish some global configurations, such as defining the number of observations (trading days) and contracts. Additionally, we select the model from which the simulated data will be generated.
Figure 4: Specify model parameters (Schwartz and Smith model).
For the Schwartz-Smith model, we assume that the logarithm of the spot price $S_{t}$ is the sum of two hidden factors:
$\log{(S_{t})}=\chi_{t}+\xi_{t},$
where $\chi_{t}$ represents the short term fluctuation and $\xi_{t}$ is the long term equilibrium price level. Both $\chi_{t}$ and $\xi_{t}$ are assumed to follow a risk-neutral mean-reverting process:
$d\chi_{t}=(-\kappa\chi_{t}-\lambda_{\chi})dt+\sigma_{\chi}dW_{t}^{\chi*},$
$d\xi_{t}=(\mu_{\xi}-\gamma\xi_{t}-\lambda_{\xi})dt+\sigma_{\xi}dW_{t}^{\xi*}.$
Here, $\kappa,\gamma\in\mathbb{R}^{+}$ are the speed of mean-reversion parameters, controlling how quickly these two latent factors converge to their mean levels. Most experiments suggest that $\kappa$ and $\gamma$ fall within the range $(0,3]$. To avoid issues with parameter identification, we recommend setting $\kappa$ greater than $\gamma$, meaning the short-term fluctuation factor converges faster than the long-term factor. The parameter $\mu_{\xi}\in\mathbb{R}$ represents the mean level of the long-term factor $\xi_{t}$, and we assume the short-term factor converges to zero. The parameters $\sigma_{\chi},\sigma_{\xi}\in\mathbb{R}^{+}$ are volatilities, representing the degree of variation in the price series over time. $\lambda_{\chi},\lambda_{\xi}\in\mathbb{R}$ are risk premiums. While we price commodities under the arbitrage-free assumption, in reality, mean term corrections, represented by $\lambda_{\chi}$ and $\lambda_{\xi}$, are necessary. $W_{t}^{\chi*}$ and $W_{t}^{\xi*}$ are correlated standard Brownian Motions with a correlation coefficient $\rho$. In discrete time, these processes are discretised to Gaussian noises, and all of these are parameters that need to be specified. Additionally, for simplicity, we assume that the standard errors $\sigma_{i},i=1,\dots,m$ for futures contracts are evenly spaced, i.e., $\sigma_{1}-\sigma_{2}=\sigma_{2}-\sigma_{3}=\dots=\sigma_{m-1}-\sigma_{m}$.
Figure 5: Download simulated data.
Finally, all the simulated data can be downloaded. Click the “Download prices” and “Download maturities” buttons to download futures prices and maturities data, respectively. Note that although the Schwartz and Smith model specifies the logarithm of the spot price, all downloaded or plotted data are in real prices - they have been exponentiated. The “Generate new data” button allows users to simulate multiple realisations from the same set of parameters. Clicking on it will generate another set of random noises, resulting in different futures prices. This button is optional if users only need one realisation. The data will update automatically when you change any parameters.
II. Implementation of the polynomial diffusion model
The procedure for simulating data from the polynomial diffusion model closely mirrors that of the Schwartz and Smith model, but it requires the specification of additional parameters.
Figure 6: Specify model parameters (polynomial diffusion model).
Both the polynomial diffusion model and the Schwartz and Smith model assume that the spot price $S_{t}$ is influenced by two latent factors, $\chi_{t}$ and $\xi_{t}$. However, while the Schwartz and Smith model assumes that the logarithm of $S_{t}$ is the sum of these two factors, the polynomial diffusion model posits that $S_{t}$ takes on a polynomial form. Currently, the PDSim GUI supports polynomials of degree 2, which can be expressed as:
$S_{t}=\alpha_{1}+\alpha_{2}\chi_{t}+\alpha_{3}\xi_{t}+\alpha_{4}\chi_{t}^{2}+\alpha_{5}\chi_{t}\xi_{t}+\alpha_{6}\xi_{t}^{2}.$
The coefficients $\alpha_{i},i=1,\dots,6$ are additional parameters specific to the polynomial diffusion model. If users wish to specify a polynomial of degree 1, they can simply set $\alpha_{4}=\alpha_{5}=\alpha_{6}=0$. Additionally, users are required to choose one of the non-linear filtering methods: the Extended Kalman Filter (EKF) or the Unscented Kalman Filter (UKF). All other procedures for simulating data follow the same steps as those in the Schwartz and Smith model.
Quality control
We conducted several tests to ensure that all functionalities of PDSim are operating correctly.
Firstly, we replicated some results from [17] using PDSim. Our replications precisely match the two figures from [17], as discussed in the section “Replicating Schwartz and Smith’s Results” on the GitHub page.
Secondly, we plotted the simulated and estimated futures prices with 95% confidence intervals for both the Schwartz and Smith model and the polynomial diffusion model. These plots specifically validate the implementation of the filtering methods. Users can reproduce these plots using the code provided in the sections “Tests for Schwartz and Smith Model” and “Tests for Polynomial Diffusion Model” on the GitHub page.
Finally, users can perform a unit test via the “Unit Tests” navigation bar of PDSim. This unit test calculates the coverage rate, defined as the proportion of simulated trajectories where over 95% of points fall within the confidence interval for a specific set of parameters. If the coverage rate exceeds 95%, we consider the set of parameters to be reasonable. A detailed introduction to the unit test is available in the “Unit Tests” section on the GitHub page.
(2) Availability
Operating system
PDSim can be run on any operating system that supports R. Additionally, it can be run on the Shiny server at https://peilunhe.shinyapps.io/pdsim/ or via Docker.
Programming language
R 4.3.0 or above.
Additional system requirements
None.
Dependencies
PDSim requires the following R packages:
  • DT ($>=0.31$)
  • ggplot2 ($>=3.4.4$)
  • lubridate ($>=1.9.3$)
  • MASS ($>=7.3.60$)
  • plotly ($>=4.10.4$)
  • scales ($>=1.3.0$)
  • shiny ($>=1.8.0$)
  • shinythemes ($>=1.2.0$)
  • tidyr ($>=1.3.1$)
List of contributors
All authors contributed to the software.
Software location:
Code repository
Name:
GitHub
Persistent identifier:
Licence:
MIT
Date published:
14/08/24
Language
English
(3) Reuse potential
Some of the functions in PDSim were originally developed and used in [10] to estimate hidden state variables and futures prices using the Schwartz-Smith model and the polynomial diffusion model. This package not only allows for detailed estimation but also offers a dynamic web application that vividly visualises the parameter identification challenges discussed in [10]. It serves as a valuable tool for researchers and practitioners in finance who are focused on commodity futures pricing. The interactive interface provides users with a deep understanding of both models, especially by illustrating the impact of individual parameter changes. The user-friendly design ensures accessibility for all users, even those with no programming experience. Additionally, the downloadable plots can be seamlessly incorporated into professional reports or academic papers.
Acknowledgements
This study was presented at the 17th International Conference on Computational and Financial Econometrics, Mathematics of Risk 2022, 4th Australasian Commodity Markets Conference, and 24th International Congress on Insurance: Mathematics and Economics. We would like to thank the audiences and organisers for their valuable feedback and suggestions.
Funding statement
This software is not funded by any individuals or institutions.
Competing interests
The authors declare that they have no competing interests.